observation and calculations. class 11 th chemistry project. plz help me GUAVA SOLUTION BURETTE READING VOLUME OF THE KMnO4 used in ml INITIAL FINAL

Grade:12

3 Answers

SAGAR SINGH - IIT DELHI
13 years ago

Dear laly,

Make the pulp of guava and boil it with dil. H2SO4.Then titrate it with standard KMnO4(in burette) solution till u get light pink colour.Repeat this with 1,2 and 3 days old guava fruit.
let weight of guava extract taken each time=50 g
Volume of guava extract each time in titration =10.0 ml
Normality of KMnO4 solution=1/20.
then
N1V1 = N2V2
(guava extract) (KMnO4 sol)
N1 * 10=1/20 *x
Normality of oxalate, N1=x/200

Please feel free to ask your queries here. We are all IITians and here to help you in your IIT JEE preparation.

All the best.

Win exciting gifts by answering the questions on Discussion Forum. So help discuss any query on askiitians forum and become an Elite Expert League askiitian.

Now you score 5+15 POINTS by uploading your Pic and Downloading the Askiitians Toolbar  respectively : Click here to download the toolbar..

 

Askiitians Expert

Sagar Singh

B.Tech, IIT Delhi

SAGAR SINGH - IIT DELHI
13 years ago

Dear laly,

I have replies to your query, please post the same ques only once...

Please feel free to ask your queries here. We are all IITians and here to help you in your IIT JEE preparation.

All the best.

Win exciting gifts by answering the questions on Discussion Forum. So help discuss any query on askiitians forum and become an Elite Expert League askiitian.

Now you score 5+15 POINTS by uploading your Pic and Downloading the Askiitians Toolbar  respectively : Click here to download the toolbar..

 

Askiitians Expert

Sagar Singh

B.Tech, IIT Delhi

SAGAR SINGH - IIT DELHI
13 years ago

Dear student,

Please do not post the same doubt again n again...

Please feel free to ask your queries here. We are all IITians and here to help you in your IIT JEE preparation.

All the best.

Win exciting gifts by answering the questions on Discussion Forum. So help discuss any query on askiitians forum and become an Elite Expert League askiitian.

Now you score 5+15 POINTS by uploading your Pic and Downloading the Askiitians Toolbar  respectively : Click here to download the toolbar..

 

Askiitians Expert

Sagar Singh

B.Tech, IIT Delhi